The perimeter of a rectangle is 256. The length is 3 times the width. What are the length and width of the rectangle?
Answer by TimothyLamb(4379)

2L + 2w = 256
L = 3w
---
2L + 2w = 256
2(3w) + 2w = 256
6w + 2w = 256
8w = 256
---
w = 32
L = 96
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php